The derivative of F(x) =12xe^6x


is


a.12e^6x (1 + x)


b.12e^6x (1+ 6x^2)


c.36e^6x (1+x)


d.12e^6x (1+6x)


1
Expert's answer
2022-04-22T05:47:18-0400

F’(x)=(g(x)*f(x))’= g’(x)*f(x)+g(x)*f’(x)

F’(x)=(12xe6x)’=12e6x+12x*6e6x=12e6x(1+6x)

answer: d